if critical angle for material of prism is C and angle of prism is A then there will be no emergent ray when
​a )A2C d) A
how explain the concept

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
From the prism properties,
we know that i1+i2=A+s
where the following parameters are defined as:
i1=angle of incidence
i2=angle of emergence
A=Angle of prism

s=Angle of deviation

Whether we want grazing emergence or grazing incidence the condition is same for both of them. i.e i1=i2

In that condition Angle of refraction =A/2

For no emergent ray A/2 >=C
A>=2C
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)